# Construct to Commitment: The Effect of Narratives on Economic Growth

### v1

@@ 1 Introduction
To bridge this gap, our NCC framework focuses on the interim implementation process, conceptualizing policy execution as a dynamic continuum of narrative dissemination and cognitive internalization. Integrating ex-ante attention to expectation formation with ex-post emphasis on causal effects, this framework reveals how policy narratives influence cognitive frameworks of market entities, thereby achieving sustainable long-term growth. The NCC framework systematically examines three dimensions: Narrative, Construct, and Commitment, respectively delineating how policy narratives systematically shape market expectations; how initial policy dissemination builds social cognitive consensus; and how narratives evolve into stable beliefs and enduring growth drivers through prolonged practice. This approach enables analysis of transmission and cognitive challenges during policy implementation while providing governments with a dynamic, process-oriented tool to enhance narrative-to-execution effectiveness.
@@ 2 Narratives-Construct-Commitment: An Overview
However, macro-level narratives alone cannot effectively guide specific market expectations and societal cognition. Thus, we introduce Construct as the critical bridge between narratives and concrete policy commitments. Unlike broad narratives, Construct emphasizes the systematic and structured refinement of policy narratives, requiring governments to translate abstract visions into specific and well-defined social expectations and cognitive frameworks through carefully designed linguistic and communicative strategies. Construct does not represent final policy commitments but rather the preliminary process of anchoring expectationstransforming abstract visions into actionable directions and psychological foundations for societal cognition. The effectiveness of Construct determines whether policy narratives can establish broad and stable belief systems within society, thereby influencing economic decisions and resource allocation among market entities.
@@
The relationships can be summarized as follows: Narrative provides the foundational framework for policy dissemination, offering macro-level visions and potential bases for societal consensus; Construct systematizes and structures narratives to shape and stabilize specific expectations and cognition among market entities, laying groundwork for policy execution; Commitment represents the final stage where policies, after long-term societal cognition and practical validation, become endogenous and institutionalized as sustainable tools for long-term economic growth.
@@ 2.2 Construct
The term construct is widely used in academia to express the meaning of social construct, emphasizing that social facts do not exist naturally but are created through language, culture, and collective understanding (Berger & Luckmann 1966). We explicitly define Construct as the initial building phase of policy narratives, a dynamic policy mechanism that systematically influences market and social cognition through narratives and discourse systems, laying the cognitive foundation for policy institutionalization and long-term economic growth (Commitment). Construct influences how market entities and the general public understand policies, thereby guiding long-term development expectations. It not only sets the direction of policies but also builds a belief in the policy vision, making it the cognitive basis for economic and social development. Construct aims to shape consensus, enhance trust, guide expectations, and is a key link in transforming policies from ideas to reality. Construct is similar to the previously mentioned Delphic Guidance, emphasizing the initial role of government policy narratives in shaping market expectations, cognitive frameworks, and investment decisions, but Construct covers a broader range of policies, not limited to monetary policy or financial markets.
@@
  • The core objective of Construct is to shape market entities’ policy expectations, forming consensus among various sectors of society regarding policy direction. It relies on Policy Narratives and Framing to transform policy objectives into an understandable, acceptable, and anticipated social cognitive system. The effectiveness of Construct depends on the systematization of policy narratives, the clarity of discourse systems, and the consistency of policy signals. The government needs to ensure the coherence and comprehensibility of policy information, transforming macro and abstract policy narratives into more specific, perceptible policy signals, enabling market entities to accurately anticipate policy directions and optimize economic behaviors such as investment, production, and consumption;
@@
In this process, Construct transforms abstract policy objectives into specific expectations and action guidelines for market entities through systematic, multi-level narrative construction, thereby producing substantive economic impacts, prompting economic entities such as enterprises, financial institutions, and local governments to actively adjust their investment and decision-making behaviors and resource allocation even before seeing specific policy measures. At this point, although the specific details of the policy have not yet been introduced, the expectations of the public and market entities have initially been anchored, and social belief in policy objectives has begun to form gradually. It can be seen that Construct, as a narrative mechanism, first establishes consensus on the importance of policies at the cognitive level among markets and the public, and initially reduces the resistance and economic uncertainty that may be encountered during future policy implementation.
@@ 2.3 Commitment
From the perspective of growth theory, Commitment is closely related to endogenous growth which emphasizes that sustained economic growth stems from technological innovation, human capital accumulation, and knowledge spillover effects within the economic system (Romer 1990; Lucas 1988). Through sustained and stable policy commitments, the government significantly reduces the uncertainty faced by enterprises in long-term innovation investment and human capital accumulation, thereby increasing the expected rate of return for economic entities and enhancing their willingness for innovation input and long-term investment (Acemoglu, Johnson, and Robinson 2005). We believe this mechanism reduces the negative impact of economic fluctuations on long-term investment decisions, enhances firms’ belief in innovation returns, and subsequently promotes corporate R&D investment and human capital input, thereby endogenously increasing the long-term productivity growth of the economy, successfully implementing policy narratives as growth commitments.
Continue with the example above, the government clearly proposes and gradually implements a long-term AI industry development strategy. After the effective dissemination of preliminary narratives in the Construct phase, the government further establishes clear fiscal policy support, legal guarantees, and long-term incentive mechanisms. For example, by setting up stable research funds, introducing industry preferential policies, improving intellectual property protection systems, and maintaining a relevant institutional environment for the long term. These measures gradually convince market entities that the development of the AI industry is not just a short-term focus of the government but a long-term and stable direction for economic growth, and policies will not fundamentally change due to short-term economic fluctuations or political changes.
@@ 3 The Model
Section I qualitatively explored the dynamic interaction process between government in Narrative, Construct, and Commitment (NCC framework), analyzing the impact of information asymmetry between central and local governments, differences in local governance capabilities, and credibility accumulation on policy implementation effects. To further clarify the internal logic of these mechanisms and determine the optimal policy strategy combination through quantitative analysis, we introduce a recursive structure model that formalizes the NCC framework as a dynamic Bayesian game between central, local governments, and market entities. In this model we treat policy narrative precision and policy implementation consistency as two cost signals released by the government (dual signals), with market entities dynamically updating their perception of government credibility based on these signals, ultimately triggering institutional commitment.
Our model is inspired by Contract Theory. Similar to Spence’s (1973) signaling model, we view the narrative precision ($\alpha$) set by the government and the implementation consistency ($\beta$) of local governments as cost signals of government type (governance capability), helping market entities identify unobservable governance capabilities through observable signals. Referencing Akerlof (1970) and Stiglitz (1975)’s screening theory, we assume that market participants use Bayesian updating mechanisms to correct their judgment of central government credibility based on local governments’ actual implementation ($\gamma$), avoiding market failure caused by information asymmetry. Drawing on Baker et al. (2016)’s policy shock identification method, we examine the different impacts of narrative shocks (short-term policy fluctuations) and implementation shocks on long-term institutional lock-in in a dynamic context, thereby clarifying how the “dual signal” mechanism promotes credibility accumulation and policy effectiveness.
@@ 3.1 Setup
We consider an infinite-horizon dynamic environment denoted as $t=0,1,2,...$. The economy consists of three types of agents: Central Government, Local Governments, and Market Participants (entities).
@@
Local Governments.Encompass multiple departments and administrative hierarchies involved in policy execution, characterized by complexity and multiplicity. For simplicity and clarity, our model synthesizes local governments into unified entities. Given disparities in economic development levels and governance capabilities, each local government faces heterogeneous execution costs represented by the intrinsic parameter $k_{c,t}$. We assume:
$k_{c,t}\sim U[\underline{k},\overline{k}],\quad 0<\underline{k}<\overline{k}<\infty.$
@@
$C_{\text{local},t}=\frac{1}{2}k_{c,t}c_{t}^{2},$
@@
$U_{\text{local},t}=\xi\theta_{t}-\frac{1}{2}k_{c,t}c_{t}^{2}-(\beta_{0}+m_{t})(c_{t}-p_{t})^{2},$
@@ 3.2 Equilibrium Analysis and Optimal Strategies
$c_{t}^{*}=\frac{(\beta_{0}+m_{t})p_{t}}{k_{c,t}+(\beta_{0}+m_{t})}.$
@@
$c_{t}^{*}=\max\left\{\frac{(\beta_{0}+m_{t})p_{t}}{k_{c,t}+(\beta_{0}+m_{t})},\,c_{\text{min}}\right\}.$
@@
${E}[U_{\text{local}}(c_{t})]=\alpha\theta_{t}-\frac{1}{2}k_{c,t}{E}[c_{t}^{2}]-(\beta_{0}+m_{t}){E}[(c_{t}-p_{t})^{2}].$
@@
$\frac{\partial{E}[U_{\text{local}}]}{\partial c_{t}^{*}}=-k_{c,t}c_{t}^{*}-2(\beta_{0}+m_{t})(c_{t}^{*}-p_{t})=0.$
@@ 4.1 Why the Innovation-Driven Development Strategy as a Case
Specifically, we regard the announcement on May 19, 2016 as the precision variable shock ($p_{t}$) of narratives in the theoretical model, and the immediate financial market responses at the time of announcements serve as the high-frequency measure of such narrative shocks($\nu_{t}$). One of the core mechanisms in our model is the dynamic updating of market participants’ evaluations of government credibility ($\theta_{t}$) during the policy implementation process. Although $\theta_{t}$ cannot be directly observed, we capture the real-time adjustment process of market expectations using financial market instantaneous responses (such as changes in stock indices within the window of policy announcements). This approach effectively maps the feedback loop of the dynamic Bayesian updating mechanism for market entities (rational participants and skeptics) in the model. Finally, when credibility accumulates to a threshold value, policies will achieve institutional commitment ($L_{t}$), entering a phase of sustained stable implementation. We have also established corresponding observation criteria in our empirical analysis, observing whether policy narrative shocks can, after generating immediate responses to market participants’ expectations (Construct phase), further continuously impact real economic indicators (such as innovation activities, investment, production, etc.) to form long-term stable trend changes. Through the impulse response functions (IRFs) of the empirical model, we can directly test the existence of this long-term commitment effect.
@@ 5.1 Narrative Shock
In our theoretical model in II, we assumed heterogeneous market participants divided according to belief processing differences (Believers, Rational Participants, Skeptics), and incorporated risk-neutral settings, thus allowing different entities to have divergent speeds and directions in response to policy shocks. However, heterogeneity in reality is often not only vertical (based on cognitive or behavioral differences) but also includes horizontal dimensions, meaning that different sectors and types of listed companies do not respond with the same intensity or manner to policy information. Therefore, we construct high-frequency narrative shocks using the following financial market variables classified by enterprise size and industry characteristics:
@@
  • 50ETF Index. This index fund tracks the constituent stocks of the SSE 50 Index, mainly covering large market capitalization blue-chip enterprises, and can more precisely reflect the immediate reactions of institutional investors and large market entities, tending toward a rational or medium to long-term pricing perspective;
@@
${28}\nu_{t}^{\text{Narrative}}=\frac{1}{3}\left(\nu_{\text{CSI300},t}+\nu_{\text{50ETF},t}+\nu_{\text{ChiNext},t}\right).$
@@ 5.2 Local Projection Redux
  • Manufacturing Fixed Asset Investment Growth: More specifically characterizes changes in enterprise-level investment behavior, representing the specific actions of enterprise entities after policy narratives and belief changes, serving as an important connecting indicator between the Construct and Commitment stages of policy;
@@ 6 Implication and Discussion
Our analysis centers on a dynamic Bayesian game framework in which Innovation-Driven Development Strategy (IDDS) announcements act as narrative shocks, shaping agents’ expectations and resource allocation. In contrast to standard “one-shot” policy interventions, the IDDS relies on a sustained feedback loop between central government signals and local implementation efforts. Under this mechanism, the credibility of the policy commitment emerges only after repeated and consistent reinforcement of the original narrative, leading to the eventual institutionalization of industrial and fiscal directives. A crucial insight is that industrial policy unfolds in a more complex ecosystem than often acknowledged by traditional models. Our findings suggest that short-run market responses to IDDS announcements (e.g., spike in technology expenditure, delayed reaction in GDP) derive not from a single exogenous shock, but from a continuous narrative constructed by central authorities and subsequently validatedor, in some cases, underminedby local enforcement and budgetary frameworks. This endogenously evolving credibility underpins our NCC (Narratives–Construct–Commitment) approach, in which initial policy articulation (Narratives) triggers expectation anchoring (Construct), ultimately culminating in stable, self-enforcing growth drivers (Commitment)  [^4] The last three sections are preliminary and we are working on improving it..
@@
Appendix A Proof of Proposition 1 (Stochastic DP Existence)
Proposition: Let the state spaces $\Theta\subseteq[0,1]$, $L\subseteq[0,L_{\max}]$, and $\Omega\subseteq[0,1-\alpha_{B}]$ be compact. Assume:
  • $\delta\in(0,1)$;
  • $U_{S}(S,p,m)$ is bounded and jointly continuous in $(p,m,S)$;
  • $f(S,p,m;\varepsilon,\nu)$ is continuous in $(\varepsilon,\nu)$, so that $\mathrm{E}[V(\cdot)]$ is well-defined;
  • the local best response $c^{*}(p,m)$ is continuous in $(p,m)$ and truncated to $[c_{\min},1]$.
Then there exists at least one Markov-perfect strategy $\{p^{*}(S),m^{*}(S)\}$ and value function $V^{*}(S)$ satisfying the stochastic Bellman equation
$V(S)=\max_{p,m}\,\mathrm{E}\bigl{[}\,U_{S}(S;p,m)\;+\;\delta\,V\bigl{(}f(S,p,m;\varepsilon,\nu)\bigr{)}\bigr{]},$
which together with $c^{*}(p,m)$ form a dynamic Stackelberg equilibrium.
Proof: Let
$\mathscr{V}=\bigl{\{}V:\mathscr{S}\to\mathbb{R}\mid V\text{ bounded, continuous}\bigr{\}},\quad\|\!V\!\|_{\infty}=\sup_{S\in\mathscr{S}}|V(S)|,\quad\mathscr{S}=\Theta\times L\times\Omega.$
Define the Bellman operator $\mathcal{T}:\mathscr{V}\to\mathscr{V}$ by
$(\mathcal{T}V)(S)=\max_{p,m}\,\mathrm{E}\bigl{[}\,U_{S}(S;p,m)\;+\;\delta\,V\bigl{(}f(S,p,m;\varepsilon,\nu)\bigr{)}\bigr{]}.$
For any $V_{1},V_{2}\in\mathscr{V}$ and all $S\in\mathscr{S}$,
$\bigl{|}(\mathcal{T}V_{1})(S)-(\mathcal{T}V_{2})(S)\bigr{|}$ $\leq\max_{p,m}\mathrm{E}\Bigl{[}\delta\bigl{|}V_{1}(f(S,p,m;\cdot))-V_{2}(f(S,p,m;\cdot))\bigr{|}\Bigr{]}$
$\leq\delta\,\|V_{1}-V_{2}\|_{\infty}.$
Since $\delta<1$, $\mathcal{T}$ is a contraction on the complete metric space $(\mathscr{V},\|\cdot\|_{\infty})$, so by Banach’s fixed‐point theorem there is a unique $V^{*}\in\mathscr{V}$ with $\mathcal{T}V^{*}=V^{*}$. Then for each $S\in\mathscr{S}$ define measurable selectors
$(p^{*}(S),m^{*}(S))=\arg\max_{p,m}\,\mathrm{E}\bigl{[}\,U_{S}(S;p,m)\;+\;\delta\,V^{*}(f(S,p,m;\varepsilon,\nu))\bigr{]}.$
Continuity and compactness ensure the maximum is attained, and measurable‐selection arguments yield Borel‐measurable strategies. Together with $c^{*}(p,m)$ these constitute the desired equilibrium.
Finally, verify the Stackelberg conditions:
  • Local Government Optimality: By Lemma 1,
    $\frac{\partial}{\partial c}\Bigl{[}\alpha\theta-\tfrac{1}{2}\kappa c^{2}-\beta(c-p)^{2}\Bigr{]}=-\kappa c-2\beta(c-p)=0\;\;\Longrightarrow\;\;c^{*}=\frac{\beta p}{\kappa+\beta}.$
  • Central Government Optimality: By construction, $(p^{*},m^{*})$ maximizes the Bellman operator.
  • Belief Consistency: Updates
    $\theta_{t+1}=\theta_{t}+\eta\,(c_{t}p_{t}-\theta_{t})+\varepsilon_{t}$
    follow Bayesian learning and remain consistent.
Some further technical points:
  • Continuity Preservation: Maximum exists by continuity of $U_{S}$ and compactness of actions.
  • Measurable Selection: Borel‐measurable selectors exist by standard theorems.
  • Admissible Strategies: Truncation $c^{*}\in[c_{\min},1]$ ensures feasibility.
∎Q.E.D.
Appendix B Constant Belief Justification
In our model, Skeptics are assumed to form a constant belief about government policy credibility, denoted by $\theta$. We specify that $\theta$ can be taken as the “historical average” of credibility. Below we justify why this choice based on the affine dependence on $\theta$.
In the aggregate investment function, $\theta$ enters linearly as part of the Skeptics’ investment term. Only the product $\omega_{t}k_{S}\theta$ appears for the Skeptics, with no nonlinear or higher-order terms in $\theta$. Hence, any constant shift of $\theta$ by a fixed amount would merely translate the level of that term.
All model equations and equilibrium conditions depend on $\theta$ only in this linear fashion. Constant shift does not alter any dynamic feedback or best-response conditions for other agents, because those rely on differences or linear combinations of credibility. Consequently, the model’s outcomes are isomorphic.
Since any constant $\theta$ is effectively just a baseline level around which the Skeptics do not update, it is convenient to interpret $\theta$ as the historical average of credibility. The linearity implies that no generality is lost by doing so.

### v3

@@ 1 Introduction
To bridge this gap, our NCC framework focuses on the interim implementation process, conceptualizing policy execution as a dynamic continuum of narrative dissemination and cognitive internalization. Integrating ex-ante attention to expectation formation with ex-post emphasis on causal effects, this framework reveals how policy narratives influence cognitive frameworks of market participants, thereby achieving sustainable long-term growth. The NCC framework systematically examines three dimensions: Narrative, Construct, and Commitment, respectively delineating how policy narratives systematically shape market expectations; how initial policy dissemination builds social cognitive consensus; and how narratives evolve into stable beliefs and enduring growth drivers through prolonged practice. This approach enables analysis of transmission and cognitive challenges during policy implementation while providing governments with a dynamic, process-oriented tool to enhance narrative-to-execution effectiveness.
@@ 2 Narratives-Construct-Commitment: An Overview
However, macro-level narratives alone cannot effectively guide specific market expectations and societal cognition. Thus, we introduce Construct as the critical bridge between narratives and concrete policy commitments. Unlike broad narratives, Construct emphasizes the systematic and structured refinement of policy narratives, requiring governments to translate abstract visions into specific and well-defined social expectations and cognitive frameworks through carefully designed linguistic and communicative strategies. Construct does not represent final policy commitments but rather the preliminary process of anchoring expectationstransforming abstract visions into actionable directions and psychological foundations for societal cognition. The effectiveness of Construct determines whether policy narratives can establish broad and stable belief systems within society, thereby influencing economic decisions and resource allocation among market participants.
@@
The relationships can be summarized as follows: Narrative provides the foundational framework for policy dissemination, offering macro-level visions and potential bases for societal consensus; Construct systematizes and structures narratives to shape and stabilize specific expectations and cognition among market participants, laying groundwork for policy execution; Commitment represents the final stage where policies, after long-term societal cognition and practical validation, become endogenous and institutionalized as sustainable tools for long-term economic growth.
@@ 2.2 Construct
The term construct is widely used in academia to express the meaning of social construct, emphasizing that social facts do not exist naturally but are created through language, culture, and collective understanding (Berger & Luckmann 1966). We explicitly define Construct as the initial building phase of policy narratives, a dynamic policy mechanism that systematically influences market and social cognition through narratives and discourse systems, laying the cognitive foundation for policy institutionalization and long-term economic growth (Commitment). Construct influences how market participants and the general public understand policies, thereby guiding long-term development expectations. It not only sets the direction of policies but also builds a belief in the policy vision, making it the cognitive basis for economic and social development. Construct aims to shape consensus, enhance trust, guide expectations, and is a key link in transforming policies from ideas to reality. Construct is similar to the previously mentioned Delphic Guidance, emphasizing the initial role of government policy narratives in shaping market expectations, cognitive frameworks, and investment decisions, but Construct covers a broader range of policies, not limited to monetary policy or financial markets.
@@
  • The core objective of Construct is to shape market participants’ policy expectations, forming consensus among various sectors of society regarding policy direction. It relies on Policy Narratives and Framing to transform policy objectives into an understandable, acceptable, and anticipated social cognitive system. The effectiveness of Construct depends on the systematization of policy narratives, the clarity of discourse systems, and the consistency of policy signals. The government needs to ensure the coherence and comprehensibility of policy information, transforming macro and abstract policy narratives into more specific, perceptible policy signals, enabling market participants to accurately anticipate policy directions and optimize economic behaviors such as investment, production, and consumption;
@@
In this process, Construct transforms abstract policy objectives into specific expectations and action guidelines for market participants through systematic, multi-level narrative construction, thereby producing substantive economic impacts, prompting economic participants such as enterprises, financial institutions, and local governments to actively adjust their investment and decision-making behaviors and resource allocation even before seeing specific policy measures. At this point, although the specific details of the policy have not yet been introduced, the expectations of the public and market participants have initially been anchored, and social belief in policy objectives has begun to form gradually. It can be seen that Construct, as a narrative mechanism, first establishes consensus on the importance of policies at the cognitive level among markets and the public, and initially reduces the resistance and economic uncertainty that may be encountered during future policy implementation.
@@ 2.3 Commitment
From the perspective of growth theory, Commitment is closely related to endogenous growth which emphasizes that sustained economic growth stems from technological innovation, human capital accumulation, and knowledge spillover effects within the economic system (Romer 1990; Lucas 1988). Through sustained and stable policy commitments, the government significantly reduces the uncertainty faced by enterprises in long-term innovation investment and human capital accumulation, thereby increasing the expected rate of return for economic participants and enhancing their willingness for innovation input and long-term investment (Acemoglu, Johnson, and Robinson 2005). We believe this mechanism reduces the negative impact of economic fluctuations on long-term investment decisions, enhances firms’ belief in innovation returns, and subsequently promotes corporate R&D investment and human capital input, thereby endogenously increasing the long-term productivity growth of the economy, successfully implementing policy narratives as growth commitments.
Continue with the example above, the government clearly proposes and gradually implements a long-term AI industry development strategy. After the effective dissemination of preliminary narratives in the Construct phase, the government further establishes clear fiscal policy support, legal guarantees, and long-term incentive mechanisms. For example, by setting up stable research funds, introducing industry preferential policies, improving intellectual property protection systems, and maintaining a relevant institutional environment for the long term. These measures gradually convince market participants that the development of the AI industry is not just a short-term focus of the government but a long-term and stable direction for economic growth, and policies will not fundamentally change due to short-term economic fluctuations or political changes.
@@ 3 The Model
Section I qualitatively explored the dynamic interaction process between government in Narrative, Construct, and Commitment (NCC framework), analyzing the impact of information asymmetry between central and local governments, differences in local governance capabilities, and credibility accumulation on policy implementation effects. To further clarify the internal logic of these mechanisms and determine the optimal policy strategy combination through quantitative analysis, we introduce a recursive structure model that formalizes the NCC framework as a dynamic Bayesian game between central, local governments, and market participants. In this model we treat policy narrative precision and policy implementation consistency as two cost signals released by the government (dual signals), with market participants dynamically updating their perception of government credibility based on these signals, ultimately triggering institutional commitment.
Our model is inspired by Contract Theory. Similar to Spence’s (1973) signaling model, we view the narrative precision ($p$) set by the government and the implementation consistency ($c$) of local governments as cost signals of government type (governance capability), helping market participants identify unobservable governance capabilities through observable signals. Referencing Akerlof (1970) and Stiglitz (1975)’s screening theory, we assume that market participants use Bayesian updating mechanisms to correct their beliefs of government credibility based on local governments’ actual implementation ($\theta$), avoiding market failure caused by information asymmetry. Drawing on Baker et al. (2016)’s policy shock identification method, we examine the different impacts of narrative shocks (short-term policy fluctuations) and implementation shocks on long-term institutional lock-in in a dynamic context, thereby clarifying how the “dual signal” mechanism promotes credibility accumulation and policy effectiveness.
@@ 3.1 Setup
We consider an infinite-horizon dynamic environment denoted as $t=0,1,2,$.... The economy consists of three types of agents: Central Government, Local Governments, and Market Participants (participants).
@@
Local Governments.Encompass multiple departments and administrative hierarchies involved in policy execution, characterized by complexity and multiplicity. For simplicity and clarity, our model synthesizes local governments into unified participants. Given disparities in economic development levels and governance capabilities, each local government faces heterogeneous execution costs represented by the intrinsic parameter $\kappa_{c,t}$. We assume:
$\kappa_{c,t}\sim U[\underline{\kappa},\overline{\kappa}],\quad 0<\underline{\kappa}<\overline{\kappa}<\infty.$
@@
$C_{\text{local},t}=\frac{1}{2}\kappa_{c,t}c_{t}^{2},$
@@
$U_{\text{local},t}=\xi\theta_{t}-\frac{1}{2}\kappa_{c,t}c_{t}^{2}-(\beta_{0}+m_{t})(c_{t}-p_{t})^{2},$
@@ 3.2 Equilibrium Analysis and Optimal Strategies
$c_{t}^{*}=\frac{(\beta_{0}+m_{t})p_{t}}{\kappa_{c,t}+(\beta_{0}+m_{t})}.$
@@
$c_{t}^{*}=\max\left\{\frac{(\beta_{0}+m_{t})p_{t}}{\kappa_{c,t}+(\beta_{0}+m_{t})},\,c_{\text{min}}\right\}.$
@@
${E}[U_{\text{local}}(c_{t})]=\alpha\theta_{t}-\frac{1}{2}\kappa_{c,t}{E}[c_{t}^{2}]-(\beta_{0}+m_{t}){E}[(c_{t}-p_{t})^{2}].$
@@
$\frac{\partial{E}[U_{\text{local}}]}{\partial c_{t}^{*}}=-\kappa_{c,t}c_{t}^{*}-2(\beta_{0}+m_{t})(c_{t}^{*}-p_{t})=0.$
@@ 4.1 Why the Innovation-Driven Development Strategy as a Case
Specifically, we regard the announcement on May 19, 2016 as the precision variable shock ($p_{t}$) of narratives in the theoretical model, and the immediate financial market responses at the time of announcements serve as the high-frequency measure of such narrative shocks($\nu_{t}$). One of the core mechanisms in our model is the dynamic updating of market participants’ evaluations of government credibility ($\theta_{t}$) during the policy implementation process. Although $\theta_{t}$ cannot be directly observed, we capture the real-time adjustment process of market expectations using financial market instantaneous responses (such as changes in stock indices within the window of policy announcements). This approach effectively maps the feedback loop of the dynamic Bayesian updating mechanism for market participants (rational participants and skeptics) in the model. Finally, when credibility accumulates to a threshold value, policies will achieve institutional commitment ($L_{t}$), entering a phase of sustained stable implementation. We have also established corresponding observation criteria in our empirical analysis, observing whether policy narrative shocks can, after generating immediate responses to market participants’ expectations (Construct phase), further continuously impact real economic indicators (such as innovation activities, investment, production, etc.) to form long-term stable trend changes. Through the impulse response functions (IRFs) of the empirical model, we can directly test the existence of this long-term commitment effect.
@@ 5.1 Narrative Shock
In our theoretical model in II, we assumed heterogeneous market participants divided according to belief processing differences (Believers, Rational Participants, Skeptics), and incorporated risk-neutral settings, thus allowing different participants to have divergent speeds and directions in response to policy shocks. However, heterogeneity in reality is often not only vertical (based on cognitive or behavioral differences) but also includes horizontal dimensions, meaning that different sectors and types of listed companies do not respond with the same intensity or manner to policy information. Therefore, we construct high-frequency narrative shocks using the following financial market variables classified by enterprise size and industry characteristics:
@@
  • 50ETF Index. This index fund tracks the constituent stocks of the SSE 50 Index, mainly covering large market capitalization blue-chip enterprises, and can more precisely reflect the immediate reactions of institutional investors and large market participants, tending toward a rational or medium to long-term pricing perspective;
@@
$\nu_{t}^{\text{Narrative}}=\frac{1}{3}\left(\nu_{\text{CSI300},t}+\nu_{\text{50ETF},t}+\nu_{\text{ChiNext},t}\right).$
@@ 5.2 Local Projection Redux
  • Manufacturing Fixed Asset Investment Growth: More specifically characterizes changes in enterprise-level investment behavior, representing the specific actions of enterprise participants after policy narratives and belief changes, serving as an important connecting indicator between the Construct and Commitment stages of policy;
@@ 6 Implication and Discussion
Our analysis centers on a dynamic Bayesian game framework in which Innovation-Driven Development Strategy (IDDS) announcements act as narrative shocks, shaping agents’ expectations and resource allocation. In contrast to standard “one-shot” policy interventions, the IDDS relies on a sustained feedback loop between central government signals and local implementation efforts. Under this mechanism, the credibility of the policy commitment emerges only after repeated and consistent reinforcement of the original narrative, leading to the eventual institutionalization of industrial and fiscal directives. A crucial insight is that industrial policy unfolds in a more complex ecosystem than often acknowledged by traditional models. Our findings suggest that short-run market responses to IDDS announcements (e.g., spike in technology expenditure, delayed reaction in GDP) derive not from a single exogenous shock, but from a continuous narrative constructed by central authorities and subsequently validatedor, in some cases, underminedby local enforcement and budgetary frameworks. This endogenously evolving credibility underpins our NCC (Narratives–Construct–Commitment) approach, in which initial policy articulation (Narratives) triggers expectation anchoring (Construct), ultimately culminating in stable, self-enforcing growth drivers (Commitment)  [^4] We are working on improving the local projection section..